OpenStudy (anonymous):

write each expression in radical notation. Evaluate the expression by hand when possible (1/125)^-1/3

jhonyy9 (jhonyy9):

5

OpenStudy (lalaly):

\[=\huge{(\frac{125}{1})^{\frac{1}{3}}}=\sqrt[3]{125}=5\]